40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5553     Entry time: Tue Sep 27 04:13:22 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: tonight's locking activity 

The lock of PRMI wasn't so robust although it could stay locked for more than 10 minutes.

There have been 2-3Hz spikes in everywhere. It needs to be investigated.

 

(to do)

 + Diagnosis on the suspensions.

 + Check the beam centering on the RFPDs.

 + Check the f2a filters on PRM and BS.

 + Health check of the suspensions by locking some cavities and measuring the noise spectra for comparison.

 + Trying to use another signal port other than AS55.

 

(Spikes)

 The attached picture below is an example of the REFLDC and POXDC signals in time series.

This was when PRCL and MICH were locked by REFL33_I and AS55_Q respectively.

PRMI.png

Note that when PRMI is unlocked, REFLDC goes to ~ 5000 counts and POXDC goes within ADC noise of ~ 1 counts.

According to the POP camera it looked like something was oscillating in the YAW direction which coincided with the spikes.

I tried finding any suspicious angular motions in the ITMs, BS and PRM olevs, but none of them showed the 2-3 Hz feature.

ELOG V3.1.3-